OpenStudy (anonymous):

I really need help. Can someone help me with a few problems? Select the equation of a line that is perpendicular to the line on the graph and passes through the point (3, 2).

OpenStudy (loser66):

Where is the graph?

OpenStudy (loser66):

Can you write out the equation of the graph?

OpenStudy (anonymous):

Is it y= -1/3 x + 2?

OpenStudy (anonymous):

Is it the right one though? :)

OpenStudy (loser66):

Nope

OpenStudy (loser66):

y = -1/3 x + not 2!!!

OpenStudy (loser66):

mistake at somewhere but it should be +3

OpenStudy (aripotta):

sorry, didn't see that it had to pass through point (3,2). loser66 is right, it should be y = -1/3x + 3

OpenStudy (anonymous):

thank you. Can you help me with one more?

OpenStudy (aripotta):

sure